# Domain adaptation, Explainability & Fairness in AI for Medical Image Analysis: Diagnosis of COVID-19 based on 3-D Chest CT-scans

### 04 Mar 2024 v1

@@ 4.1 COVID-19 detection & domain adaptation baselines
The baseline architecture adopted for both Challenges, namely the COVID-19 Detection Challenge and the Covid-19 Domain Adaptation Challenge, is a CNN-RNN architecture [5, 10, 11, 12, 13, 14, 15, 16, 17, 18, 4, 10, 19, 6, 20, 8, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 12, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@ 4.2 Pre-Processing & Implementation Details
In the pre-processing stage, all 2-D CT slices have been extracted from respective DICOM images. Next, voxel intensity values were computed through a window of $350$ Hounsfield units (HU)/$-1150$ HU; they were then normalized in the range $[0,1]$. Data augmentation was also performed, including random rotation in [-10°, 10°] and horizontal flip [46, 47, 48] to extract region of interests, such as lung areas in the 2-D images.

### 10 Mar 2024 v2

@@ 4.1 COVID-19 detection & domain adaptation baselines
The baseline architecture adopted for both Challenges, namely the COVID-19 Detection Challenge and the Covid-19 Domain Adaptation Challenge, is a CNN-RNN architecture [5, 10, 11, 12].
@@ 4.2 Pre-Processing & Implementation Details
In the pre-processing stage, all 2-D CT slices have been extracted from respective DICOM images. Next, voxel intensity values were computed through a window of $350$ Hounsfield units (HU)/$-1150$ HU; they were then normalized in the range $[0,1]$. Data augmentation was also performed, including random rotation in [-10°, 10°] and horizontal flip [14, 15] to extract region of interests, such as lung areas in the 2-D images.